# Fine-Grained Reward Optimization for Machine Translation using Error Severity Mappings

### v2

@@ 5.2 Results and Main Findings
For all language directions reported in Table 2, “tRL w/ xCOMET-MQM“ outperforms SFT and its sentence-level counterpart “sRL w/ xCOMET-MQM“ across all neural metrics considered. SFT significantly improves translation quality by tailoring the pre-trained MT model to the specific target language pairs. Moreover, applying RL methods (sRL or tRL) on top of SFT further enhances the MT model’s performance by directly optimizing translations based on targeted reward signals. When comparing the sRL and tRL methods, we observe that sRL leads to moderate improvements over SFT, while the gains obtained by tRL are more substantial, particularly when assessed with advanced neural metrics. Although the ChrF scores for tRL models trained with xCOMET-MQM are lower, possibly due to the focus on optimizing token-level features over character-level accuracy, all other metrics indicate significant improvements, suggesting the overall translation quality is better.
@@
Our severity-based fine-grained reward optimization mechanism significantly improves translation quality across all automatic evaluation metrics as shown in Table 3. These findings highlight that tRL not only improves the quality of state-of-the-art MT models but can also significantly boost stronger MT-LLMs, demonstrating its broad applicability and potential for advancing multilingual translation systems.
@@
The role of the reward model in achieving alignment is crucial, as also evidenced by our findings (Tables 2 and 3). Our results clearly show that using xCOMET as a reward model, particularly at the token level, significantly improves translation quality as measured by several metrics. Given that xCOMET exhibits a strong correlation with human judgments, it proves to be an essential tool for guiding MT models toward higher translation quality. In contrast, traditional metrics like BLEU, based on -gram overlap, can fall short in aligning with human judgments as they do not capture contextual nuances and semantic understanding Freitag et al. (2022). Consequently, BLEU performs less effectively compared to neural metrics like xCOMET in this setup which use contextual embeddings. Therefore, incorporating neural metrics as reward models is crucial for capturing the subtleties of language and improving the overall quality and reliability of MT models.
@@ 5.3 Ablation Study
We investigate the impact of different severity maps on token-level RL training using xCOMET, as detailed in Table 4. The severity maps we evaluate include the default MQM-based map (MQM), our custom map (Our), the reversed MQM-based map (rMQM), the reversed custom map (rOur), and a binary map (Bin). Our findings, shown in Table 5, highlight the importance of having gradual transitions between reward values. Smooth transition severity map to result in better translation quality. In contrast, abrupt changes in the reward signal can destabilize learning, leading to inconsistent training, oscillations, or convergence to suboptimal policies Ranzato et al. (2015); Sutton and Barto (2018). Additionally, we find that the binary severity map, which ignores the severity of errors, provides less informative feedback to the model, resulting in slightly lower performance than maps that offer more nuanced assessments. Although designing custom severity maps can increase complexity and require hyperparameter tuning, our experiments suggest they can be set in a straightforward way with minimal overhead. We leave a more systematic investigation of these mappings, including the possibility of learning them during training, to future work.
@@
Figure 3 shows the evolution of mean rewards during training for sRL and tRL methods across the two datasets for the NLLB system. As can be observed, tRL training shows a more stable and consistently increasing reward evolution, which is crucial for the training process as it ensures consistent improvements and reduces the risk of overfitting or fluctuations that could degrade the MT model’s performance.
tRL improves translation quality for longer sequences.
Building on our hypothesis that tRL is particularly effective for longer sequences, we show the COMET22 scores for the WMT24 EN DE dataset, grouped by sequence length, across different training methods, including NLLB and Tower, in Figure 4. tRL consistently outperforms other methods, especially for longer sequences, which we attribute to its ability to capture localized and fine-grained signals. These results further support our findings, showing that tRL remains the most robust method, with scores dropping less than other approaches as the source sentence length increases. This reinforces the effectiveness of tRL in handling complex, longer sentence structures.
@@
We compare REINFORCE and PPO for RL-based MT with xCOMET as reward model, evaluating their impact on translation quality (Table 6). Both methods are effective, but PPO achieves superior overall metric scores due to features such as objective clipping and KL divergence control, which enhance training stability. However, REINFORCE remains a strong alternative for simpler implementations that aim to achieve competitive performance.
| Reward Model | BLEU | xCOMET |
| Latency | ||
| Throughput | ||
| Quality |
@@
tRL using xCOMET offers higher quality, as measured by COMET22, than tRL with BLEU but comes with increased computational costs, including higher latency and lower throughput due to its larger pre-trained encoder, as reported in Table 7. Despite the slower processing speeds, xCOMET’s superior reward quality and resulting higher model quality are particularly valuable in token-level feedback scenarios, where high-quality rewards are essential, and the additional computational cost is manageable with GPU acceleration. Furthermore, advances in compute efficiency for transformer-based architectures, such as quantization Jacob et al. (2017), FlashAttention Dao et al. (2022); Dao (2023), and distillation Hinton et al. (2015), can help reduce the computational load, making xCOMET more practical for broader applications.
@@ 5.4 Human Evaluation
For our human evaluation, we used Direct Assessments (Graham et al., 2013, DAs) to score translations on a scale from 0 to 100, following the standard WMT human evaluation methodology. We evaluate 200 randomly chosen instances from the WMT18 EN DE dataset. Two professional translators, both native speakers of the target language, assesse the references and NLLB translations using the following methods: SFT, CPO with xCOMET, tRL with xCOMET and our proposed severity mapping, and sRL with xCOMET.
While WMT18 EN DE allows for easy comparison between several methods due to shorter sequences, we conducte a second human evaluation on the WMT24 EN DE dataset to validate our empirical finding that tRL benefits longer input sequences. For this setting, we directly compare sRL with xCOMET and tRL with xCOMET in a pairwise setting using outputs from the Tower model. To ensure adequate coverage across different sequence lengths, we performed stratified sampling based on source length, using bins [0, 100, 250, 500, 1000] with 50 instances per bin.
| Model | Ann1 | Ann2 |
| CPO | 59.5 | 53.2 |
| SFT | 64.5 | 54.2 |
| tRL | 66.9 | 56.1 |
| sRL | 66.8 | 57.6 |
| Reference | 76.0 | 60.0 |
@@
As shown in Table 8, both sRL and tRL models consistently outperform SFT and CPO, demonstrating their advantage in enhancing translation quality. On average, human evaluation scores for tRL and sRL are similar. On the WMT24 dataset, tRL achieves an average DA score of 82.6, 2.3 points higher than sRL. We additionally show the distribution of DA scores bucketed by length in Figure 5. Across all buckets, tRL outperforms sRL, with more significant (based on error bars overlap) gains for longer sentences. These findings are also consistent with our automatic evaluation, including the ablation analysis (§ 5.3) and clearly demonstrate that tRL enhances stability and translation quality, particularly for longer sentences.

### v3

@@ 5.2 Results and Main Findings
For all translation directions reported in Table 2, “tRL w/ xCOMET-MQM“ outperforms SFT and its sentence-level counterpart “sRL w/ xCOMET-MQM“ across all neural metrics considered. SFT significantly improves translation quality by tailoring the pre-trained MT model to the specific target language pairs. Moreover, applying RL methods (sRL or tRL) on top of SFT further enhances the MT model’s performance by directly optimizing translations based on targeted reward signals. When comparing the sRL and tRL methods, we observe that sRL leads to moderate improvements over SFT, while the gains obtained by tRL are more substantial, particularly when assessed with advanced neural metrics. Although the chrF scores for tRL models trained with xCOMET-MQM are lower on IWSLT2017, this is likely due to a mismatch between the reward signal and the evaluation metric: token-level rewards optimize semantic quality, not character-level overlap. This can lead to fluent, accurate translations that diverge lexically from references, reducing chrF despite improved overall quality. Neural metrics, which are more robust to surface-level variation and better aligned with human judgments freitag-etal-2022-results; freitag-etal-2023-results, as well as human evaluation (see Section 5.4) – the gold standard for assessing translation quality – consistently show improvements with our tRL approach. Appendix C presents a focused quantitative and qualitative analysis of cases in which chrF decreases while xCOMET improves.
@@
Our severity-based, fine-grained mechanism significantly improves translation quality across all automatic evaluation metrics, as shown in Table 3. These findings highlight that tRL not only improves the quality of state-of-the-art MT models but can also significantly boost stronger LLM-based MT systems, demonstrating its broad applicability and potential for advancing multilingual MT systems.
@@
The role of the reward model in achieving alignment is crucial, as also evidenced by our findings (Tables 2 and 3). Our results clearly show that using xCOMET as a reward model, particularly at the token level, significantly improves translation quality as measured by several metrics. Given that xCOMET exhibits a strong correlation with human judgments, it proves to be an essential tool for guiding MT models toward higher translation quality. In contrast, traditional metrics like BLEU, based on -gram overlap, can fall short in aligning with human judgments as they do not capture contextual nuances and semantic understanding freitag-etal-2022-results. Consequently, BLEU performs less effectively compared to neural metrics like xCOMET in this setup which use contextual embeddings. Therefore, incorporating neural metrics as reward models is crucial for capturing the subtleties of language and improving the overall quality and reliability of MT models.
@@ 5.3 Ablation Study
We investigate the impact of different severity maps on token-level RL training using xCOMET, as detailed in Table 4. The severity maps we evaluate include the default MQM-based map (MQM), our custom map (Our), the reversed MQM-based map (rMQM), the reversed custom map (rOur), and a binary map (Bin). Our findings, shown in Table 5, highlight the importance of having gradual transitions between reward values. Smooth transition severity map to result in better translation quality. In contrast, abrupt changes in the reward signal can destabilize learning, leading to inconsistent training, oscillations, or convergence to suboptimal policies RANZATO2016; sutton2018reinforcement. Additionally, we find that the binary severity map, which ignores the severity of errors, provides less informative feedback to the model, resulting in slightly lower performance than maps that offer more nuanced assessments. Although designing custom severity maps can increase complexity and require hyperparameter tuning, our experiments suggest they can be set in a straightforward way with minimal overhead. We leave a more systematic investigation of these mappings, including the possibility of learning them during training, to future work.
@@
Figure 3 shows the evolution of mean rewards during training for sRL and tRL across the two datasets for the NLLB system. As observed, tRL training exhibits a more stable and consistently increasing reward trajectory, which is crucial for ensuring steady improvements and reducing the risk of performance-degrading fluctuations or overfitting.
@@
tRL improves translation quality for longer sequences.
Building on our hypothesis that tRL is particularly effective for longer sequences, we present COMET22 scores for the WMT24 ENDE dataset in Figure 4, grouped by source sequence length and comparing different training methods, including NLLB and Tower. The figure also shows the distribution of source sequence lengths in the training and test data. Notably, the training data is skewed toward shorter inputs—a common characteristic of large MT corpora—whereas the WMT24 test set includes a broader distribution with a higher proportion of long sequences. This discrepancy highlights the need for models that generalize well to longer inputs. In this context, "longer sequences" refers to short paragraphs rather than single sentences. These are not document-level inputs and remain within the 512-token limit of xCOMET, ensuring that the reward model processes them without truncation during training. tRL consistently outperforms other training methods, especially on longer sequences. We attribute this to its ability to capture localized, fine-grained reward signals during training. These findings further support our earlier results: tRL demonstrates the most robust performance, with smaller performance drops as the source sentence length increases, confirming its strength in handling complex sentence structures. For completeness, we also evaluate in Appendix B a hybrid model that applies sRL to short inputs and tRL to long ones. This approach outperforms sRL but remains inferior to tRL.
@@
We compare REINFORCE and PPO for RL-based MT with xCOMET as the reward model, evaluating their impact on translation quality (Table 6). Both methods are effective, but PPO achieves superior overall metric scores due to features such as objective clipping and KL divergence control, which enhance training stability. However, REINFORCE remains a strong alternative for simpler implementations that aim to achieve competitive performance.
@@
Using xCOMET as a reward model for tRL yields higher-quality translations than BLEU, but it also incurs increased computational costs, as detailed in Table 7. Due to its larger pre-trained encoder, xCOMET exhibits significantly higher latency (average seconds per token) and lower throughput (tokens per second). It is worth noting that throughput values can appear higher than latency alone might suggest, as this metric benefits from amortized per-call overheads and batching. The quality improvement, measured by COMET22, reflects the performance of our best model (Tower) on WMT24 ENDE after fine-grained optimization with each reward model. Despite the slower processing, xCOMET’s superior reward quality is particularly valuable in token-level feedback scenarios where high-quality rewards are essential and the additional computational cost is manageable with GPU acceleration. Furthermore, advances in computational efficiency for transformer architectures—such as quantization quant, FlashAttention fa; fa2, and distillation dist—can help mitigate this computational load, making xCOMET more practical for broader applications.
| Latency | Throughput | Quality | |
| BLEU | |||
| xCOMET |
@@ 5.4 Human Evaluation
For our human evaluation, we used Direct Assessments (graham-etal-2013-continuous, DAs) to score translations on a scale from 0 to 100, following the standard WMT human evaluation methodology. We evaluate 200 randomly chosen instances from the WMT18 EN DE dataset. Two professional translators, both native speakers of the target language, assess the references and NLLB translations using the following methods: SFT, CPO with xCOMET, tRL with xCOMET, our proposed severity mapping, and sRL with xCOMET.
While WMT18 EN DE allows for easy comparison between several methods due to shorter sequences, we conducted a second human evaluation on the WMT24 EN DE dataset to validate our empirical finding that tRL benefits longer input sequences. For this setting, we directly compare sRL with xCOMET and tRL with xCOMET in a pairwise setting using outputs from the Tower model. To ensure adequate coverage across different sequence lengths, we performed stratified sampling based on source length, using bins [0, 100, 250, 500, 1000] with 50 instances per bin.
| CPO | SFT | tRL | sRL | Reference | |
| Ann1 | 59.5 | 64.5 | 66.9 | 66.8 | 76.0 |
| Ann2 | 53.2 | 54.2 | 56.1 | 57.6 | 60.0 |
@@
As shown in Table 8, both sRL and tRL models consistently outperform SFT and CPO, demonstrating their advantage in translation quality. On the WMT24 dataset, tRL achieves an average DA score of , points higher than sRL, with consistent gains across sentence-length buckets, as shown in Figure 5; longer sentences exhibit more significant improvements based on error bar overlap. Human evaluations were conducted by two professional annotators, and inter-annotator agreement is moderate-to-high (Pearson’s , Spearman’s ), indicating reliable scoring. These results align with our automatic evaluation, including the ablation analysis, which collectively shows that tRL enhances stability and translation quality, particularly for longer sentences.
@@ Acknowledgments
Appendix A Details of the Severity Assignment Algorithm
In this section, we provide a detailed description of the severity assignment algorithm used in our approach, focusing on the case of tokens overlapping annotated error spans. Our implementation resolves multiple overlapping error spans by assigning the token the worst severity among all overlapping spans. This choice aligns with MQM annotation practices, where the most severe issue in a region governs its quality classification.
Formally, if a token overlaps spans labeled minor, major, and critical, the token is assigned critical. We avoid averaging or length-weighted schemes to remain fully tokenizer-agnostic. A token is considered affected by a span if it overlaps with it in any way, not only if it is fully contained. This avoids mismatches in cases where tokens are longer than the annotated spans.
Formally, let
be the set of tokens,
be the set of annotated error spans,
denote the severity of span .
The severity assigned to token is defined as:
Example.
Suppose the text is tokenized as a single token “abc”, and error spans are defined as (minor), (major), and (critical). Since the token “abc” overlaps with all three spans, its assigned severity is critical.
Appendix B Details of the Hybrid sRL–tRL Experiment
| Method | Metrics | |||||
| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | |
| Tower | 42.77 | 62.60 | 71.80 | 88.50 | 68.59 | 67.20 |
| + SFT | 45.14 | 63.30 | 72.18 | 88.90 | 69.18 | 67.30 |
| + sRL w/ xCOMET-MQM | 45.58 | 63.76 | 72.41 | 90.20 | 70.55 | 69.70 |
| + tRL w/ xCOMET-MQM | 46.92 | 65.63 | 74.66 | 91.90 | 71.80 | 71.20 |
| + hRL w/ xCOMET-MQM | 45.96 | 65.07 | 73.20 | 90.57 | 71.30 | 69.34 |
To examine whether the improvements of tRL are primarily driven by long-sequence behavior, we evaluate a hybrid reinforcement learning (hRL) approach that applies sentence-level RL (sRL) to short inputs and token-level RL (tRL) to long ones. Short sentences are defined as those below the average source length in the training data, and long sentences as those above it. This experiment follows our best-performing configuration: the Tower model with xCOMET-MQM used as the reward signal. Table 9 reports the results alongside the relevant baselines from Table 3. The hRL model improves over sRL on most metrics but consistently falls short of tRL, indicating that tRL provides the strongest training signal across sentence lengths without the added complexity of a hybrid setup.
Appendix C Analysis of chrF Drop Cases
This section provides a focused quantitative and qualitative analysis of translations where chrF decreases but xCOMET improves under token-level RL (tRL). The goal is to examine whether the observed chrF drop corresponds to genuine translation degradation or reflects a metric mismatch. We leverage available human evaluation data for WMT24 ENDE, using Direct Assessment (DA) scores as reliable indicators of translation quality.
This analysis stems from the observation that, in some settings, tRL outputs show a noticeable drop in chrF (up to 7 points on IWSLT2017) while simultaneously improving xCOMET and other quality metrics. This raised concerns that such a drop might reflect lexical imprecision or other undesirable artifacts. To investigate, we conducted a detailed error analysis focusing on cases with the largest discrepancies between chrF and xCOMET. Examining these high-divergence examples provides the clearest insight into whether chrF drops correspond to real quality issues or simply reflect a mismatch between metrics.
C.1 Quantitative Analysis
Figure 6 summarizes the extreme discrepancy cases for ENDE translations. The plot shows the average DA score for the top-N cases with the largest chrF–xCOMET discrepancies. Even with the observed chrF drops and corresponding xCOMET increases, tRL consistently achieves higher human DA scores than sRL, indicating that translation quality is not compromised. This strongly suggests that the apparent chrF decline is a metric artifact rather than a genuine degradation in translation quality.
C.2 Qualitative Analysis
Table 10 presents representative examples of translations with chrF drops but higher xCOMET and DA scores. These cases illustrate how lexical divergence from the reference can lower chrF while yielding translations that are more fluent, semantically accurate, and preferred by human evaluators.
| Example 1 | |
| Source | I like both the glasses you posted, these are a really nice colour…makes me want new ones |
| Reference | Ich mag beide Brillen, die du gepostet hast, die haben eine wirklich nette Farbe … weckt bei mir den Wunsch, neue zu kaufen |
| sRL translation | Ich mag beide Brillen, die du gepostet hast, diese sind wirklich eine schöne Farbe… Ich würde gerne neue haben |
| tRL translation | Mir gefallen beide Brillen, die du gepostet hast, sie haben eine wirklich schöne Farbe… jetzt will ich auch neue |
| Scores | ChrF 63.4 59.7 xCOMET: 96.8 99.3 DA: 75.0 84.0 |
| Example 2 | |
| Source | I don’t have any telemetry for the battery. This is something I might add in a future revision. |
| Reference | Ich habe keine Telemetrie für den Akku. So etwas könnte ich in Zukunft noch einbauen. |
| sRL translation | Ich habe keine Telemetrie für den Akku. Das ist etwas, was ich in einer zukünftigen Überarbeitung hinzufügen könnte. |
| tRL translation | Ich habe keine Telemetrie-Daten für den Akku. Das ist etwas, was ich vielleicht in einer zukünftigen Version hinzufügen könnte. |
| Scores | ChrF 58.4 53.1 xCOMET: 97.2 98.2 DA: 84.0 94.0 |
| Example 3 | |
| Source | “Which notebook is that?” Ivory asked, sitting down next to Kari on her bed. |
| Reference | „Welches Notizbuch ist das?“, fragte Ivory und setzte sich neben Kari auf deren Bett. |
| sRL translation | „Welches Notizbuch ist das?“ fragte Ivory und setzte sich neben Kari aufs Bett. |
| tRL translation | „Welches Notizbuch ist das?“, fragte Ivory, während sie sich neben Kari auf das Bett setzte. |
| Scores | ChrF 85.3 80.4 xCOMET: 99.3 99.5 DA: 84.0 93.0 |
| Example 4 | |
| Source | I’m splurging on a new set of frames, these red ones I reeeeally like. |
| Reference | Ich verschwende mein Geld für eine neue Brillenfassung, diese rote mag ich seeehr. |
| sRL translation | Ich habe mir eine neue Brille gekauft, diese rote gefällt mir richtig gut. |
| tRL translation | Ich gönne mir ein neues Brillengestell, diese roten gefallen mir ganz besonders. |
| Scores | ChrF 37.8 33.8 xCOMET: 95.2 95.5 DA: 76.0 92.0 |
| Example 5 | |
| Source | “Well my buddys from boot camp here did” Cohren motioned to Harris, Craith and Ravik. “But then again, their squad and platoon leader’s, they know to wake up before their units. |
| Reference | „Nun, bei meinem Kumpels aus dem Bootcamp ist das bereits der Fall“, Cohren winkte Harris, Craith und Ravik zu. „Aber auch ihre Truppen- und Zugführer wissen, dass sie vor ihren Einheiten aufwachen müssen.“ |
| sRL translation | „Nun ja, meine Kameraden vom Bootcamp hier haben es getan“, sagte Cohren zu Harris, Craith und Ravik. „Aber ihr Zug- und Platoonführer wissen, dass sie vor ihren Einheiten aufwachen müssen. |
| tRL translation | „Nun, meine Kameraden vom Bootcamp haben es getan“, sagte Cohren und deutete dabei auf Harris, Craith und Ravik. „Aber sie und ihre Zugführer wissen ja auch, dass sie vor ihren Einheiten aufstehen müssen.“ |
| Scores | ChrF 57.8 56.2 xCOMET: 81.9 85.1 DA: 74.0 85.0 |